OpenStudy (zeesbrat3):

The surface area, S, of a sphere of radius r feet is S = S(r) = 4πr2. Find the instantaneous rate of change of the surface area with respect to the radius r at r = 8.

OpenStudy (zeesbrat3):

So do you find r'? Which would end up as 64π?

ganeshie8 (ganeshie8):

Careful, you're finding S'(r) here and yes it ends up being 64pi

OpenStudy (plasmataco):

yes, I think so

OpenStudy (zeesbrat3):

You mean I might be learning something?! Finally!! Thank you everyone

OpenStudy (anonymous):

the instantaneous rate of change of the surface area\[ \frac{dS}{dr} \]

OpenStudy (anonymous):

\[ \frac{dS}{dr} = \frac{d\left(4\pi r^2\right)}{dr} \]

OpenStudy (zeesbrat3):

Then the product rule if I remember the name correctly.. Right?

OpenStudy (anonymous):

You can use the product rule, but the power rule is intersting.

OpenStudy (zeesbrat3):

Sorry, I meant the power rule. Where you multiply the value of the exponent by the coefficient and then decrease the exponent by 1

OpenStudy (anonymous):

See what happens

OpenStudy (zeesbrat3):

8πr 8π(8) 64π

OpenStudy (anonymous):

Remember units

OpenStudy (zeesbrat3):

ft^3 right?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

When you substituted in \(r=8\), you should have put in the \(\text {ft}\) as well.